5.35*1021 formula units of sodium sulfate, Na2SO4 

Express your answer to three significant figures and include the appropriate units. 
Answer #1

use:

number of mol = number of molecules / Avogadro’s number

number of mol = 5.35*10^21 / 6.022*10^23 mol

number of mol = 8.884*10^-3 mol

Molar mass of Na2SO4,

MM = 2*MM(Na) + 1*MM(S) + 4*MM(O)

= 2*22.99 + 1*32.07 + 4*16.0

= 142.05 g/mol

use:

mass of Na2SO4,

m = number of mol * molar mass

= 8.884*10^-3 mol * 1.421*10^2 g/mol

= 1.262 g

Answer: 1.26 g